A 65 kg mountain climber climbs through a vertical rope at a constant acceleration from rest to a




1 of 1.25 m/s in 10.0 s and feels no appreciable air resistance. How much power must the




speed m/s in 10.0s and alr must




climber imparts on the vertical rope?

1
Expert's answer
2022-04-18T17:31:54-0400

The acceleration of the climber is:


"a = \dfrac{v_f}{t} = \dfrac{1.25m/s}{10s} = 0.125m/s^2"

where "v_f = 1.25m\/s" is its final speed, and "\Delta t = 10s" is the time. The velocity is then depends on time as follows:


"v(t) = at"


According to the second Newton's law, the net force on the climber is:


"F = ma"

where "m = 65kg" is his mass.

The power is then:


"P(t) = F\cdot v(t) = ma\cdot at = ma^2t"

This expression depends on time. I assume it is asked to find the average power during 10s. The average power is:


"P_{av} = \dfrac{1}{\Delta t}\int_0^{\Delta t} P(t)dt = \dfrac{1}{\Delta t}\int_0^{\Delta t} ma^2tdt=ma^2\dfrac{\Delta t}{2}\\
P_{av} = \dfrac{65kg\cdot (0.125m/s^2)^2\cdot 10s}{2} \approx 5.1W"

Answer. 5.1W.